# Mamo: a Mathematical Modeling Benchmark with Solvers

### v2

@@ 5.4 Data Statistics
In the section on Ordinary Differential Equations (ODE), we present a total of 346 problems: 196 are based on first-order equations, 110 on second-order equations, and 40 on systems of equations, offering a comprehensive exploration across different levels. The optimization section is divided into two segments: Easy LP, featuring 652 high school-level Mixed Integer Linear Programming (MILP) problems, and Complex LP, comprising 211 undergraduate-level problems that integrate both LP and MILP. See the word cloud of the combined data in the Appendix L.
@@ 6.1.1 Evaluation Details
Our evaluation protocol is meticulously designed to test the LLMs’ mathematical modeling process. For ODE problems, the LLM is prompted to generate Python code from a natural language description. The code is executed, and its output is compared with the correct answer to assess accuracy. In optimization problems, the LLM is asked to express the model in .lp format, after which the COPT solver is used to find the optimal value for comparison with the correct answer.
@@
In ODE parts, we ask LLMs to output Python code calling solvers ( solve_ivp odeint in SciPy, and dsolve in NumPy). In the optimization sections, we prompt LLMs to generate the standard .lp file. Compared to the testing methodology in NLP4LP (AhmadiTeshnizi et al., 2023b), which involves prompting the model to output Python code for invoking the solver, the .lp format more closely aligns with the standard optimization form, making it more readable. Furthermore, optimization solvers such as COPT (Ge et al., 2023) can directly read and solve the .lp file, minimizing potential issues related to the LLM’s coding capabilities. To better adhere to the format, we conducted tests using 3-shot learning. Detailed information about the number of few-shot samples used in our experiments and the result can be found in the Appendix (see Appendix I and Section O). Besides, at the beginning of the testing, we do the first clean of the codes, clearing up the string like ‘ “‘ python “‘ ’ that will affect our testing.
Mitigation of Formatting Errors
It is essential to distinguish errors in modeling from those in coding or file formatting [^3] Appendix  H displays the example of formating errors instead of modeling errors.. Therefore, if the LLM’s output contains syntax errors in python code or .lp code, we need to rectify these issues without altering the underlying logic of the code. This step ensures that our evaluation focuses solely on the modeling process. To achieve this, we use a language model to fix any syntax error, referring to it as the code modifier. In our experiments, we use the original language model and GPT-4-0613 as code modifiers. For the sake of reproducibility, we choose the original LLM; To ensure the comparison focuses exclusively on modeling capacity, we choose GPT-4-0613 as the code modifier due to its exceptional coding proficiency and operational reliability, as well as its stability than other provisional release Language Models. See the comparison of the result from different code modifiers in Section 6.3, and prompts in Appendix  O.
@@ 6.1.2 Evaluation Metrics
Given the potential for minor discrepancies in numerical solutions, our protocol includes a comparison rule that accommodates for slight inaccuracies. Let $O$ denote the output from the LLM, and $A$ denote the standard answer. The comparison is defined by the following criteria  444 See the evaluation script in Appendix M
General Comparison Criterion Let $n$ denotes the decimal places of $A$ ($n=0$ if $A$ is integer). The Answer $O$ is correct if
$\left|\frac{O-A}{A}\right|<10^{-4}\text{ or }|O-A|<\min\{10^{-n},10^{-2}\}$ (1)
This metric is designed [^5] See the case study in Appendix D to account for the precision errors inherent in solvers and the specified significant figures in questions, ensuring that the LLM’s output is accurately evaluated against the standard answer with allowances for numerical precision and variability in computational responses.
@@ 6.2 Benchmarking Results
Table 3: Evaluation Results on the Mamo Benchmark. The subscript denotes improvement using the original LLM as a code modifier, which corrects syntax errors without changing the underlying logic, differentiating modeling from formatting errors (see Section 6.1.1). The "Overall" score represents the weighted average of correct rates, weighted by question count. The highest score in each category is marked in boldface; an underline signifies the top score with the LLM itself as code modifier. The score after using LLM as code modifier is computed as the sum of the original score and the value in the subscript.
@@
First Second System Easy Complex
order (%) order (%) (%) LP (%) LP (%)
GPT-4o 67.86 +0.51 36.36 +0.91 40.00
@@
Llama-3-8b 0.51
@@
    Here GPT-4-turbo denotes GPT-4-turbo-2024-04-09; GPT-4o denotes GPT-4o-2024-05-13; GPT-4 refers to GPT-4-0613; GPT-3.5-turbo denotes GPT-3.5-turbo-0125
Our evaluation  [^6] See the setting in Appendix N includes both closed-source LLMs and open-source LLMs. The closed-source commercial LLMs include the GPT-4 series OpenAI (2023), the Claude series  [^7] https://www.anthropic.com/news/claude-3-family, and the Gemini series  [^8] https://deepmind.google/technologies/gemini/. The open-source LLMs include DeepSeek series by  DeepSeek-AI (2024),  Guo et al. (2024) and  Shao et al. (2024); Llama-3 models  [^9] https://github.com/meta-llama/llama3; Qwen-1.5 (Bai et al.) and Mixtral (Jiang et al.). Table 3 presents the evaluation results of different language models on the Mamo Benchmark. The performance of these models is assessed both in their original form and when enhanced with the language model itself to rectify syntax errors (the improvement is indicated by the subscript). Table 5 in Appendix A shows results when GPT-4-0613 is selected as the code modifier.
The evaluation results on the Mamo Benchmark reveal that GPT-4o stands out among the models in both ODE and LP tasks. Specifically, GPT-4o achieves the highest performance in ODE tasks with 68.37% in first-order and 37.27% in second-order equations, and also excels in LP tasks with 87.42% in easy LP and 31.28% in complex LP scenarios when using itself as code modifier. Additionally, DeepSeek-v2, as an open-source LLM, demonstrates significant capability, achieving a 57.57% success rate across all tasks when enhanced with syntax correction by itself, underscoring its potential despite its open-source nature. Moreover, Mixtral-8x22B also exhibits competitive performance relative to DeepSeek-v2. After fix syntax error by itself, it achieves a remarkable score of 21.33% on the complex LP task, showcasing its potential in modeling complex optimization problems.
The inclusion of the LLM itself as code modifier to rectify syntax errors results in significant improvements across various models. For instance, GPT-3.5-turbo’s performance in ODE tasks increased from 12.72% to 18.21% overall and Claude-3-Haiku shows a remarkable improvement in ODE tasks from 32.95% to 38.73%. Similarly, DeepSeek-v2, as an open-source LLM, demonstrates significant capability with its overall ODE performance increasing from 32.08% to 35.55% when enhanced with syntax correction.
6.3 On the Format Errors and Code Modifiers
@@
DeepSeek-v2 70.81 83.82 96.24 90.27 92.00 94.90
DeepSeek-coder 88.44 93.64 95.95 92.58 96.29 98.84
@@
Llama-3-8B 20.23 26.01 68.21 75.09 76.36 79.14
Llama-3-70B 24.56 30.92 77.46 77.64 79.49 89.46
Qwen-1.5-72B-Chat 55.49 89.60 95.38 77.40 92.24 95.37
Qwen-1.5-110B-Chat 51.45 63.29 95.09 71.26 84.01 90.15
@@
The data presented in Table 4 provides a comparative analysis of the performance (rate of correct format) of various models using different modification methods. The table considers three scenarios: no modification, self-modification, and modification provided by GPT-4-0613.
The GPT-3.5-turbo model shows the most dramatic improvement when modified by GPT-4, jumping from a 37.28% success rate to 93.93%. Interestingly, Claude-3 models (Sonnet and Haiku) and DeepSeek models (DeepSeek-v2, DeepSeek-coder) also exhibit high improvement rates when enhanced by GPT-4-0613, highlighting the compatibility of GPT-4’s modification methods with other models. In the realm of Optimization tasks, the GPT-3.5-turbo model again shows a significant improvement when enhanced by GPT-4, jumping from a 87.95% success rate to 96.76%.
The modification from GPT-4 is generally superior to both self-modification and no modification. This indicates that GPT-4 can serve as a more effective and fair code modifier. By using GPT-4 for syntax error correction, we can better examine the modeling ability of different LLMs without the confounding factor of formatting errors. Table 5 in Appendix A shows the results of selecting GPT-4-0613 as code modifier.
6.4 Analysis
Given the results in Section 6.26.3 and Appendix A, several observations emerge from the benchmark results. For open-source models, larger models generally perform better, as evidenced by comparisons between Llama-3-70B and Llama-3-8B, Mixtral-8x22B-instruct and Mixtral-8x7B-instruct, with the larger models consistently achieving higher scores.
From Table 4, it is apparent that some models struggle with adhering to the syntax of the solver, such as GPT-3.5-turbo and certain open-source models. Some models like Llama-3-8B may not closely follow the instructions, even worse than the syntax error.These models tend to show significant improvements in both correct rates (as indicated by the subscript in Table 3) and execution success rates after the application of code modifiers, particularly when their initial performance without modifiers is low. Notably, in the ODE part, GPT-3.5-turbo achieves a remarkable 28.03% increase in execution success rate by self-modified, accompanied by substantial improvements in ODE performance. In Optimization part, Qwen-1.5-72B-Chat follows the same trends. When comparing the results from Table 3 and Table 5, it becomes clear that the improvements are even more pronounced for open-source models when utilizing GPT-4 modifications rather than self-modifications. See Appendix E for the analysis of typical modeling errors and case study,
@@ 7 Conclusion and Future Directions
In conclusion, Mamo’s innovative benchmark shifts the evaluation focus of LLMs towards the modeling process, offering a nuanced assessment of their mathematical problem-solving capabilities. By isolating modeling from solving through the use of solvers, Mamo provides a comprehensive understanding of LLMs’ strengths and limitations in mathematical modeling. This approach not only enhances our insights into the fundamental modeling abilities of LLMs but also guides future research by highlighting the importance of the modeling process in solving mathematical problems.
@@ Appendix A GPT-4-0613 as Code Modifier
DeepSeek-v2 †† 54.08 30.91 30.00 86.81 3.79 60.05
DeepSeek-coder †† 30.10 11.82 15.00 78.83 18.96 52.27
@@
Llama-3-8b †† 11.22 0.91 0.00 63.19 8.53 37.47
Llama-3-70b †† 14.80 7.27 10.00 70.71 9.48 43.18
Qwen-1.5-72b-Chat †† 16.84 13.01 5.45 78.53 6.64 47.80
Qwen-1.5-110b-Chat †† 40.82 13.64 20.00 78.83 14.22 53.51
@@
Selecting GPT-4 as a code modifier generally results in better performance than using the LLM itself for self-improvement. For instance, DeepSeek-v2’s accuracy in ODE tasks improves from 32.08% with self-modification to 43.93% when GPT-4 is used as the code modifier. This pattern indicates that GPT-4 is a more effective code modifier, enhancing the accuracy and performance of various models more reliably than self-improvement.

### v3

@@ 3.5 Data Statistics
In the section on Ordinary Differential Equations, we present a total of 346 problems: 196 are based on first-order equations, 110 on second-order equations, and 40 on systems of equations, offering a comprehensive exploration across different levels. The optimization section is divided into two segments: Easy_LP, featuring 652 high school-level Mixed Integer Linear Programming (MILP) problems, and Complex_LP, comprising 211 undergraduate-level problems that integrate both LP and MILP. See the category statistics and the word cloud of the combined data in Appendix O.
@@ 4.1.1 Evaluation Details
Our evaluation protocol follows the methodology outlined in Section 2.3. First, the LLMs were prompted to generate formalized code (in each domain) to represent a mathematical model, denoted as $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\widehat{M}$. This code was then executed by a solver, and the result, $\widehat{A}$, was compared with the ground truth answer, $A$.
For ODE problems, the LLM was prompted to generate Python code based on a natural language description. The generated code was executed, and the output was compared to the correct solution to assess its accuracy. For optimization problems, the LLM was asked to express the model in .lp format, after which an optimization solver COPT (Ge et al., 2023) was employed to find the optimal value for comparison with the correct answer. This process is independent of the specific solver used; for instance, MATLAB (Inc., 2022) can be used as an ODE solver, and Gurobi (Gurobi Optimization, LLC, 2024) can serve as an alternative optimization solver.
Benchmarking Models
Our evaluation [^2] See the settings in Appendix Q includes both proprietary and open-source LLMs. The proprietary LLMs consist of the GPT-4 series (OpenAI, 2023), the Claude series [^3] https://www.anthropic.com/news/claude-3-family, and the Gemini series [^4] https://deepmind.google/technologies/gemini/. The open-source LLMs include the DeepSeek series (Shao et al., 2024; DeepSeek-AI, 2024), Llama-3.1 models (Dubey et al., 2024), Qwen-2.5 (Team, 2024), and Mixtral (Jiang et al., 2024).
@@
We prompted the LLMs to generate Python code that called solvers (such as solve_ivp and odeint in SciPy, and dsolve in NumPy). In the optimization sections, the LLMs were asked to generate a standard .lp file. Compared to the testing methodology in NLP4LP (AhmadiTeshnizi et al., 2023), which involved prompting the model to output Python code for invoking the optimization solver, the .lp format more closely aligned with the standard optimization form, making it more readable. Optimization solvers like COPT (Ge et al., 2023) could directly read and solve the .lp file. To better adhere to this format, we conducted tests using 3-shot learning. Detailed information about the number of few-shot samples used in our experiments and the results can be found in Appendix K and Appendix R. Additionally, at the start of the testing process, we performed an initial cleaning of the code, removing strings such as ``` python ``` that would have affected our testing.
@@ 4.1.2 Evaluation Metrics
According to Property 2 and Property 3, both the ground truth answer $A$ and the answer $\widehat{A}$, derived from solving the LLM’s output mathematical model $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\widehat{M}$, are numerical values with specified significant figures. To account for minor discrepancies, we apply the following comparison rule:
General Comparison Criterion
: Let $n$ denote the number of decimal places in $A$ ($n=0$ if $A$ is an integer). The answer $\widehat{A}$ is considered correct if:
$\left|\frac{\widehat{A}-A}{A}\right|<10^{-4}\text{ or }|\widehat{A}-A|<\min\{10^{-n},10^{-2}\}.$
Figure 4: Examples illustrating the application of the General Comparison Criterion
This rule accounts for precision errors and ensures an accurate evaluation of $\widehat{A}$ while allowing for slight computational variations. Figure 4 demonstrates the application of this criterion with examples, illustrating how discrepancies between the correct and generated answers are assessed. For further details, see Appendix P and the case study in Appendix E.
@@ 4.2 Benchmarking Results
First
Second
System
Easy
Complex
order (%)
order (%)
LP (%)
LP (%)
Proprietary Models
o1-preview 71.43 +0.00 50.00 +0.00 45.00 +0.00 80.21
+0.00
36.02 +3.79 67.60
+0.66
GPT-4o 67.86
+0.51
36.36
+0.91
40.00
@@
Llama-3.1-8B-instruct 30.61
+1.53
4.55
+1.81
10.00
+2.50
74.39
+2.60
13.27
+0.00
48.14
+1.90
Llama-3.1-70B-instruct 62.24
+0.52
33.64
+5.45
32.50
+0.00
85.74
+0.36
22.75
+1.42
64.44
+1.02
Llama-3.1-405B-instruct 66.84
+1.02
40.00
+0.00
30.00
+0.00
86.20
+0.76
34.12
+4.74
67.91 +1.40
Qwen-2.5-32B-instruct 60.71
@@
    Here GPT-4-turbo denotes GPT-4-turbo-2024-04-09; GPT-4o denotes GPT-4o-2024-05-13; GPT-4 refers to GPT-4-0613; GPT-3.5-turbo denotes GPT-3.5-turbo-0125.
Table 4: Evaluation Results on the Mamo Benchmark. The subscript denotes improvement using the original LLM as a code modifier, which corrects syntax errors without changing the underlying logic, differentiating modeling from formatting errors (see Section 4.3). The "Overall" score represents the weighted average of correct rates, weighted by question count. The highest score in each category is marked in boldface; an underline signifies the top score with the LLM itself as code modifier. The score after using LLM as code modifier is computed as the sum of the original score and the value in the subscript.
@@
Open-source Models
DeepSeek-v2.5 97.69 99.13 99.13 90.36 93.83 97.55
@@
Llama-3.1-8b-instruct 57.51 72.97 92.20 80.99 88.30 94.55
Llama-3.1-70b-instruct 92.48 98.27 97.98 92.93 96.99 97.34
Llama-3.1-405b-instruct 98.84 99.71 99.71 92.58 98.03 95.48
Qwen-2.5-32B-instruct 97.69 98.84 99.13 93.17 96.29 94.67
Qwen-2.5-72B-instruct 97.40 97.69 97.98 90.80 94.06 91.15
@@
Table 4 presented the evaluation results (accuracy rate) of different language models on the Mamo Benchmark. The performance of these models was assessed both in their original form and when enhanced by the language model (i.e., code modifier) itself to rectify syntax errors (with improvements indicated by the subscript). Appendix A provided results when GPT-4-0613 was used as the code modifier, and Appendix F presented the error analysis.
Mamo is Challenging
As the results in Table 4 indicate, even state-of-the-art models struggled with complex tasks such as second-order ODEs, systems of ODEs, and Complex_LP. This is due to the increasing number of variables and the more intricate relationships among these variables compared to easier tasks like Easy_LP.
Take-away 1.
Existing LLMs struggle with complex tasks in mathematical modeling.
The Performance of o1
Interestingly, o1-preview, although it excelled in complex tasks such as second-order ODEs and Complex_LP, did not outperform (and was even worse than GPT-3.5-Turbo) in Easy_LP tasks, which consist of high-school-level linear programming modeling problems. As shown in Table 6, the execution success rates of o1-preview in each category were nearly 100% across all methods (raw, self-modified, or GPT-4-modified), indicating that this underperformance was not due to limitations in coding ability.
Take-away 2.
o1 excelled in complex tasks but underperformed in easy tasks.
Figure 5: The testing pipeline involving code modifier.
Open-source Models vs. Proprietary Models
One of the most notable aspects of this benchmark is the emergence of open-source models as serious competitors to closed-source models. The Llama-3.1-405B, in particular, achieved the best overall score across both evaluation categories. The Llama-3.1 series, Qwen series, and DeepSeek-v2.5, despite being open-source models, demonstrated remarkable performance, even surpassing highly regarded commercially developed models such as GPT-4 and Claude-3 in categories like second-order ODEs and Optimization.
Take-away 3.
Open-source models are competitive in simpler tasks, but there is still a gap compared to closed-source models in more complex tasks.
Effectiveness of Scaling
The evaluation demonstrates a correlation between model size and performance across almost all tested categories. As indicated by the results of the Llama-3.1, Qwen, and Mixtral series, for models with the same architecture but different parameter scales (e.g., Llama-3.1-7B vs. Llama-3.1-405B), larger models tended to outperform their smaller counterparts. This observation highlights the effectiveness of scaling in the mathematical reasoning tasks of LLMs. However, the effectiveness of scaling may plateau on easier tasks, such as Easy_LP problems.
Take-away 4.
Larger LLMs perform better in mathematical modeling.
4.3 On the Impact of Format Errors
Mitigation of Formatting Errors
It is important to differentiate between errors in modeling and those related to coding or file formatting (See Appendix I). When the LLM output contains syntax errors in Python or .lp code, these errors must be corrected without altering the logic of the code to ensure that the evaluation focuses solely on modeling. To address syntax issues, we employ a code modifier, as the pipeline shown in Figure 5 (More details are discussed in Appendix D). In our experiments, we used the original language model and GPT-4-0613 as code modifiers, with the former selected for reproducibility and the latter for its superior coding proficiency and stability. A comparison of results from different code modifiers is provided in Section 4.2 and Appendix A, with prompts detailed in Appendix R.
Setting
@@ 4.3 On the Impact of Format Errors
The testing pipeline is shown in Figure 5. As mentioned in Section 4.1.1, we selected the original LLM and GPT-4-0613 as code modifiers. The data presented in Table 5 provides a comparison of the performance (rate of correct format) of various models using different modification methods: i) no modification, ii) self-modification, and iii) modification provided by GPT-4-0613.
Observation
Combining Table 4 and Table 5, the improvement can be attributed to misjudgments caused by the LLM’s coding ability. For LLMs with a lower initial correct format rate (such as GPT-3.5-Turbo and Llama-3.1-8B in ODE tasks), the relative improvement is higher. This is because, in several cases, their output codes, initially judged as "Error" due to compile errors, actually represent correct mathematical models. It is also worth noting that the improvement has little effect on the ranking, indicating that the influence of different coding abilities on our benchmark is minimal.
Take-away 5.
Format errors have minimal impact on mathematical modeling.
5 Related Work
Recent studies have advanced the use of large language models (LLMs) in mathematical problem-solving by introducing new datasets and solver methodologies. For example, Frieder et al. (2024) and Yuan et al. (2023) developed datasets to test LLMs on various mathematical problems and arithmetic expressions. Lewkowycz et al. (2022) focused on training models using natural language paired with LaTeX-formatted math from arXiv. To address complex computational tasks, Zhang et al. (2024) introduced the CARP dataset, while He et al. (2024) presented OlympiadBench, a bilingual benchmark for competition-level math reasoning. Liu et al. (2024) also developed MathBench, structured to assess LLMs’ theoretical knowledge. Additionally, Mirzadeh et al. (2024) proposed GSM-Symbolic, a benchmark that generates symbolic variants of GSM8K questions. Their study also introduced GSM-NoOp, a dataset containing irrelevant but seemingly related information to assess model robustness, demonstrating that while LLMs exhibit sensitivity to numerical variations, they remain resilient to superficial modifications.
In solver development, efforts have focused on integrating LLMs into novel problem-solving approaches. Yang et al. (2023) explored LLMs as direct solvers, highlighting their autonomous problem-solving potential. He-Yueya et al. (2023) introduced a hybrid method that combines LLMs with external symbolic solvers for equation solving, while Pan et al. (2023) developed LOGIC-LM, a framework blending LLMs with symbolic solvers to tackle logical problems. AhmadiTeshnizi et al. (2023) contributed to both dataset development and solver integration by creating NLP4LP, a benchmark for LP and MILP problems, and OptiMUS, an LLM-based agent for optimization problem solving. In line with these developments, Lyu et al. (2023) introduced Faithful-CoT, a two-stage framework that translates natural language into symbolic reasoning chains (represented as code) before solving the problem using corresponding solvers. This approach aligns closely with our benchmark’s objective of assessing the faithful translation of problems into structured mathematical formulations.
@@ 6 Conclusion and Future Directions
Mathematical modeling is a crucial yet challenging aspect of mathematical reasoning. Evaluating LLMs’ ability to handle modeling is essential, as their performance will continue to improve over time, making specialized benchmarks in specific areas necessary to accurately assess and push their capabilities. Due to the complexity of mathematical models, assessment is difficult. To address this, we introduce the Mamo benchmark, which incorporates solvers to validate answers and assess the correctness of the modeling process. Mamo demonstrates that LLMs can be effectively evaluated on complex reasoning tasks, where the correct answer may vary but remains equivalent in principle. This approach paves the way for further research into LLMs’ reasoning capabilities.
Limitations
While the current benchmarking methodology necessitates LLMs to generate Python code for ODEs and .lp files for optimization problems, it opens up exciting avenues for future research. The process, while effective, may inadvertently influence the LLMs’ focus towards formalization over the conceptual modeling, potentially affecting the depth of mathematical reasoning. This opens the door for benchmarks that better distinguish between an LLM’s modeling skills and formalization abilities. Future work could focus on assessing LLMs’ mathematical modeling directly, bypassing formalization for a more nuanced evaluation.
Ethics Statement:
Acknowledgment
This work was supported by the Shenzhen Science and Technology Program (JCYJ20220818103001002), Shenzhen Doctoral Startup Funding (RCBS20221008093330065), Tianyuan Fund for Mathematics of National Natural Science Foundation of China (NSFC) (12326608), Shenzhen Key Laboratory of Cross-Modal Cognitive Computing (grant number ZDSYS20230626091302006), and Shenzhen Stability Science Program 2023.
@@ Appendix A GPT-4-0613 as Code Modifier
Open-source Models
DeepSeek-v2.5 †† 67.86 36.36 32.50 86.20 13.33 64.20
@@
Llama-3.1-8b-instruct †† 33.67 8.18 15.00 82.06 13.27 53.27
Llama-3.1-70b-instruct †† 62.76 39.09 32.50 86.81 23.70 65.76
Llama-3.1-405b-instruct †† 67.86 40.00 30.00 86.66 34.60 68.40
Qwen-2.5-32b-instruct †† 61.73 38.18 35.00 84.05 13.27 62.28
Qwen-2.5-72b-instruct †† 60.71 37.27 32.50 89.57 11.43 64.61
@@
[…]
Selecting GPT-4 as a code modifier generally results in better performance than using the LLM itself for self-improvement. For instance, DeepSeek-v2’s accuracy in ODE tasks improves from 36.41% with self-modification to 48.80% when GPT-4 is used as the code modifier. This pattern indicates that GPT-4 is a more effective code modifier, enhancing the accuracy and performance of various models more reliably than self-improvement.